Find the slope and y-intercept of :
y=-6x+4

Answer by ramkikk66(644)


This is equation is of the form y = m*x + c, where the slope is m and y intercept
is c.

Comparing with y = -6*x + 4, we get the answer as:

slope = -6
y intercept = 4 (i.e. the point (0,4))

:)